OpenStudy (anonymous):

What is the simplified form of x/4 + y/3 over x/3 - y/4 ??

OpenStudy (callisto):

\[\frac{\frac{x}{4} +\frac{y}{3}}{\frac{x}{3} -\frac{y}{4}}\]\[=\frac{\frac{3x}{(3)(4)} +\frac{4y}{(4)(3)}}{\frac{4x}{(4)(3)} -\frac{3y}{(3)(4)}}\]\[=\frac{\frac{3x+4y}{12}}{\frac{4x-3y}{12}}\]\[=\frac{3x+4y}{12}\times\frac{12}{4x-3y}\]\[=\frac{3x+4y}{4x-3y}\]

OpenStudy (lgbasallote):

awww she latex so fast >.<

OpenStudy (callisto):

Do you understand my workings?